Fermi National Laboratory
 

Operation Summary for September 13, 2010 - September 15, 2010

 

Monday Sept. 13

The day shift began with Operations monitoring the Tevatron (TeV) store 8082 (88.02E30) and the Recycler stash (226.1E10), with the Antiproton Source (Pbar) stacking (15.4E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with MiniBooNE off for an access, with the NuMI experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, TeV personnel began conducting end-of-store studies at 10:10 AM. The MiniBooNE access ended and operators reestablished beam to the experiment at 11 AM. Operations terminated store 8082 at 12:03 PM. TeV personnel conducted a period of machine studies. Operations put the TeV into shot setup at 3:08 PM. A MI RF station (MIRF13) bias power supply developed a LCW leak; technicians installed another power supply. Interlock techs completed the Meson safety system tests.

During the evening shift, Operations established store 8085 at 4:12 PM with an average initial luminosity (AIL) of 301.9E30. Operations reported some downtime when the MI power supply interlock loop dropped out and the MI-30 VCB opened; operators quickly closed the VCB and normal beam operations resumed. Operations sent beam to the Switchyard (SY) dump.

Tuesday Sept. 14

The midnight shift began with Operations monitoring store 8085 (104.67E30) and the stash (194.9E10), with Pbar stacking (26.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8085 at 7:13 AM and put the TeV into shot setup.

During the day shift, Operations established store 8086 at 8:18 AM with an AIL of 325.7E30. At 9:21 AM, Operations reported that a longitudinal instability in the TeV caused a drop in luminosity of 20E30. Operations reported a few minutes of stacking downtime due to a bad target airflow calculation. There was also some downtime due to Linac RF trips and due to P1 line vacuum valve closures when operators moved the extraction septa into the beam.

During the evening shift, Operations reported that TRF3 tripped off due to an RF gap envelope fault; this caused a slight loss in luminosity. MI personnel conducted MI-8 line collimator studies. The MI power supply interlock loop dropped out when the VCB for the MI-30 quadrupole power supply opened; operators closed the VCB.

Wednesday Sept. 15

The midnight shift began with Operations monitoring store 8086 and the stash (356.2E10), with Pbar stacking (26.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8086 at 12:05 AM and put the TeV into shot setup, but MI SBD problems kept operators from loading protons into the TeV; Operations called in Controls personnel to investigate. CDF reported that they had a blown fuse for a crate power supply in their collision hall that required an access to change. D0 also needed a collision hall access to work on a bad power supply; operators prepared both areas for access. Operations reported that TRF3 tripped off several times during this shift; Operations called in RF techs and they reported that they were replacing the station’s power amplifier. Near the end of the shift, many Booster RF stations began sparking and tripping off; operators traced the problem to BRF5. They turned the station off, which resolved the problem.

The Plans for Wednesday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/15/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory